OpenStudy (anonymous):

Find the slope of the line through the pair of points. A(2, –3), P(2, 9)

OpenStudy (anonymous):

the slope of a line through two given points \(\large (x_1,y_1) \) and \(\large (x_2,y_2) \) is the formula: \(\large slope=m=\frac{y_2-y_1}{x_2-x_1} \)

OpenStudy (anonymous):

okay they give me these choices A. 0 B. 2/3 C. - 1/3 D. undefined

OpenStudy (anonymous):

let's calculate the values... \(\large y_2-y_1= \) and \(\large x_2-x_1= \)

OpenStudy (anonymous):

ok i just started these so i honestly dont even know

OpenStudy (anonymous):

\(\large y_2-y_1= \) means subtract the y-coordinates...

OpenStudy (anonymous):

so 1?

OpenStudy (anonymous):

\(\large x_2-x_1= \) means subtract the x-coordinates

OpenStudy (anonymous):

so it's - 1/3

OpenStudy (anonymous):

\(\large (x_1,y_1)=(2, -3) \) \(\large (x_2,y_2)=(2, 9) \)

OpenStudy (anonymous):

okay so what i am understanding from what you just wrote out is it is undefined

OpenStudy (anonymous):

\(\large (y_2-y_1)=9-(-3)= \)

OpenStudy (anonymous):

...

OpenStudy (anonymous):

i'm trying to make you figure this out if you'd only answer my questions...

OpenStudy (anonymous):

9 - (-3) = 6?

OpenStudy (anonymous):

no... 9-(-3)=12

OpenStudy (anonymous):

oh it goes up? ok that was a help next

OpenStudy (anonymous):

\(\large (x_2-x_1)= \) ???

OpenStudy (anonymous):

2 -9 ? 7

OpenStudy (anonymous):

what are the x-coordinates the the two points?

OpenStudy (anonymous):

it says A(2, –3), P(2, 9)

OpenStudy (anonymous):

so \(\large x_1=2 \) and \(\large x_2=2 \) so what's \(\large x_2-x_1= \) ???

OpenStudy (anonymous):

\(\large x_2-x_1=2-2=0 \) so \(\large slope=m=\frac{y_2-y_1}{x_2-x_1}=\frac{9-(-3)}{2-2}=\frac{12}{0}=???? \)

OpenStudy (anonymous):

12